ganeshie8 (ganeshie8):

After a job is finished on a multicore machine, I see below time/memory usage info. may i know how to interpret the terms real, user, sys, elapsed time, lvheap, malloc, cputime : --- GRAND TOTAL CPU TIME = 3509 REAL TIME = 1256 LVHEAP = 189/2492/2629 MALLOC = 3081/3081/3081 ELAPSED TIME = 1256 real 1256.44 user 3525.82 sys 31.31

thomaster (thomaster):

Real = wall clock time (from start to finish) User = amount of CPU time spend in user-mode code within the process. Sys = amount of CPU time used in executing the process. malloc = memory allocation lvheap = a pool of memory that is used in allocating portions of memory (from the heap) for requests. (not entirely sure about this one)

ganeshie8 (ganeshie8):

real time = 1256 sec = 1256/60 mnts = 21 mnts so after starting this job, il need to wait 21 mnts for it to finish is it.. malloc = 3081/3081/3081 does it mean, the job took min/avg/max of ~3GB ?

ganeshie8 (ganeshie8):

so i need to check oly for real time in the log, to get an idea of how much time it took to complete ?

thomaster (thomaster):

yea real is all elapsed time including time slices used by other processes and time the process spends blocked

ganeshie8 (ganeshie8):

got it thank you:) can we also knw the total RAM used by the job from above info ?